# NESTful: A Benchmark for Evaluating LLMs on Nested Sequences of API Calls

### v2

@@ 4.2 Experimental Settings
The experiments are carried out in one-shot and three-shot settings, i.e., the prompt contains one or three in-context learning (ICL) examples, respectively. To the best of our knowledge, all the 15 models were not trained with the label assignment syntax in the output API sequence, so it was crucial to have ICL examples to get the best results from the models. For each model, we used its specified prompt along with the special tags. Context length limitations prevented the inclusion of the entire API library in the prompt. Instead, we pre-processed the data to create a shorter API list for each sample. This list ensured the inclusion of the gold APIs, the APIs used in the in-context learning (ICL) examples, and some random APIs, keeping the total prompt length under 4k tokens (based on granite-3.0-8b-instruct tokenizer). Also, the API calls were extracted from the model’s response as a list of JSON objects, taking into account the specific prompt format and output structure for each model. Finally, we evaluated a zero-shot ReAct Yao et al. (2022) agent with the best 4 models. The Think-Act-Observe loop in ReAct was limited to 10 iterations. All experiments have employed greedy sampling and set the temperature to its lowest value of 0.0 for all the models.
@@ 4.3 Metrics
Generated response from the model is a sequence of API calls, with each call consisting of an API name and its argument-value pairs. We use the Partial Sequence Matching metric to determine how many predicted APIs (with their argument-value pairs) in a sequence match with the gold API sequence. In contrast, the Full Sequence Matching metric evaluates whether the model predicts the exact full sequence of APIs, including both the API names and their argument-value pairs, when compared to the gold API sequence. In both cases, we calculate the scores for each sample and then compute the statistical mean across the entire dataset as the final score. The final metric we use is the Win Rate, which measures if all the predicted APIs by the model are valid and when executed lead to the gold answer. Unlike the F1 and Accuracy measures, which assess the alignment of predictions with the gold API sequence, the Win Rate focuses on the final gold answer. A win is recorded if the predicted answer matches the gold answer.
Model Total Parameters Direct Prompting (One-shot ICL) ReAct Agent (Zero-shot)
F1 Func. F1 Param. Part. Acc. Full Acc. Win Rate F1 Func. F1 Param. Part. Acc. Full Acc. Win Rate
MadeAgents/Hammer2.0-7b 7B 0.62 0.42 0.29 0.25 0.33 0.56 0.34 0.14 0.00 0.25
Granite-20B-FunctionCalling 20B 0.60 0.37 0.28 0.24 0.26 0.40 0.00 0.00 0.00 0.11
llama-3-1-70b-instruct 70B 0.35 0.18 0.10 0.00 0.01 0.30 0.36 0.13 0.00 0.42
Mixtral-8x22B-Instruct-v0.1 141B 0.43 0.30 0.22 0.16 0.10 0.31 0.38 0.17 0.02 0.52
Table 2: Evaluation results on NESTful comparing the performance of a ReAct Agent (zero-shot) to the Direct Prompting with a one-shot ICL example. The best performance is highlighted in bold; the second best is underlined.
@@ 4.4 Results
Table 1 presents a comparison of different baselines on the NESTful dataset with one-shot and three-shot ICL example settings. The low numbers of the best function calling models in Table 1 depicts the complexity and toughness of the nested sequencing problem. Hammer2.0-7b achieved the highest performance with a full accuracy of 25% and a win rate of 34%, significantly below acceptable numbers for real world applications. These numbers are similar to multi-turn setting in BFCL v3 dataset. [^10] https://gorilla.cs.berkeley.edu/leaderboard.html This clearly depicts the significant scope for improvements for the models in various aspects of function calling including nested sequencing. We inspected the models’ outputs and identified several common issues across them. These models struggle with tasks such as assigning variables, utilizing output parameter details from the API specifications, and correctly passing variable names and corresponding output parameters to subsequent APIs, even with ICL examples [^11] Note: While we acknowledge that the models aren’t trained with the robust data schema discussed in Section 3.1, common issues with nested sequencing immaterial of the data schema has to improve in LLMs..
As anticipated, in most of our experiments in Table 1, the models are doing better across all the metrics when they are provided with three ICL examples in the prompt instead of one example. Across all models, Partial Sequence Match scores are consistently higher than and Full Sequence Match scores, which is expected, as the latter is a stricter metric than the former. In some cases, the Win Rate is a little higher than the Full Accuracy because models sometimes follow alternate trajectories to reach the final answer. Figure 6 shows this case where a correct final answer can be obtained by following two different solution paths. Despite the smaller model sizes, the function calling models such as Hammer2.0-7b, Granite-20B-FunctionCalling, and xLAM series did exceptionally well compared to the larger language models. Across all the metrics and evaluation settings (one shot and three shots), Hammer2.0-7B achieved either the highest or the second-highest scores. On the contrary, being a function calling model, ToolACE-8B still suffered in following the output JSON format with the proper variable assignments and that led to very bad accuracy and win-rate scores. For this benchmark, we found that the size or complexity of the model itself becomes less relevant, as long as it can effectively follow the given instructions and the ICL examples. As evidence, we can see that some larger models such as xLAM-8x22b-fc-r and llama-3-1-405b-instruct-fp8 performed poorly, whereas a 7B model (Hammer2.0-7b) was doing best.
Figure 6: An example where the predicted API sequence is valid but does not match the ground truth sequence. The final answer match results in a Win Rate of 1.
@@
Table 2 summarizes the results of the ReAct agent Yao et al. (2023) in comparison to the one-shot ICL direct prompting setting. Note that there was no ICL example provided in the ReAct case as the expected output does not need to follow the label assignment syntax. The reported metrics (except Win Rate) which measure the conformance to the ground truth API sequence show that ReAct performed poorly. This could be due to the single-step planning and execution approach as opposed to planning the entire API sequence at once.
ReAct does outperform direct prompting on Win Rate but has room for improvement. Inspecting the outputs showed that the top two models had better format alignment to ReAct and sometimes performed the functionality of some of the APIs using its parametric knowledge. This led to lower F1 and accuracy scores but higher Win Rate; the computations were correct despite mismatching steps, as shown in Figure 7.
Figure 7: An example where a ReAct agent using llama-3-1-70b-instruct used its parametric knowledge to perform a computation correctly as highlighted. This led to a final answer that matches the ground truth answer but the API sequence does not match with the ground truth sequence.
@@ 4.5 Dataset Analysis
Figure 8: Model performances with a varying number of (a) longest data dependencies, (b) the total number of data dependencies, and (c) common data nesting patterns in NESTful. We observe that models perform well for the simple data samples but the performance drops sharply with complicated structural patterns in the data.
@@
In Figure 8(b), we present the win rate compared to the total number of data dependencies within a sample (a representation of the complexity of the sequence). The trends here are similar to the ones we observed in the Nesting Depth Analysis section. Models perform well for singular nested data dependency in the sample, achieving  60% Win Rate, but the performance drops sharply with two or more nested data dependencies.
@@ 5 Challenges
Data-type Adherence for the Input/Output Parameters   In the API specification, we define the data type for all parameters— input and output. The ‘type’ field specifies the data type, such as string, number, list, etc. Since APIs follow a strict structure for both input and output, it is crucial for the model to adhere to these specified format. If the model fails to follow this, especially in cases involving nested functions where the output of one API is used as the input for another, the process will fail if the output type doesn’t match the expected input type. Additionally, in nested scenarios, the output parameters of the referenced function must be used correctly; otherwise, incorrect parameter assignments will result in an incorrect answer.
@@
In the above example, the first two functions need to be the same with different parameter values. However, for the third function subtraction, it is crucial to distinguish the outputs of the first two functions appropriately to obtain the correct result. This adds complexity to the dataset, as the models are not trained with this schema and must follow the instructions precisely. Our qualitative studies (discussed in Section 4.4) suggest the same as well.
Implicit API calling   Implicit function calling refers to a scenario where the system must invoke a specific API, along with potentially other APIs, to fulfill a user query, even though the query does not explicitly mention the task that requires that particular API. Figure 1 demonstrates an example of implicit function calling, where the user query presents an arithmetic problem without explicitly stating the APIs involved. In this case, the model must understand the problem, identify the relevant formulas, choose the appropriate function calls in sequence based on those formulas, and then populate the parameter values by either extracting information from the query or referencing the output of previous function calls. This increases the dataset complexity and makes it harder for the models.
@@ 6 Conclusion
In this work, we introduced NESTful, a new benchmark for evaluating the LLMs on nested sequences of API function calling. Existing LLMs perform poorly on this dataset as compared to their performance on existing benchmarks and identified their several modes of failure. In addition, we outlined the many challenges this dataset poses to LLM function calling approaches. By making this dataset available publicly under a permissive open-source license, we aim to push API function calling capabilities in new directions and unlock solutions to more realistic and challenging tasks.

### v3

@@ 4.2 Experimental Settings
The experiments are carried out with temperature 0.0 in one-shot and three-shot settings, i.e., the prompt contains one or three in-context learning (ICL) examples, respectively. To the best of our knowledge, all 18 open models were not trained with the label assignment syntax in the output API sequence, so it was crucial to have ICL examples to get the best results. For each model, we used its specified prompt along with the special tags. Context length limitations prevented the inclusion of the entire API library in the prompt. Instead, we pre-processed the data to create a shorter API list for each sample. This list ensured the inclusion of the gold APIs, the APIs used in the in-context learning (ICL) examples, and some random APIs, keeping the total prompt length under 4K tokens. Output API calls were extracted from the model’s response as a list of JSON objects, taking into account the specific prompt format and output structure for each model. Finally, we evaluated a zero-shot ReAct Yao et al. (2022) agent with the best 4 open models based on the win-rate and AgentLM-13B, limiting max steps to 10.
@@ 4.3 Metrics
LLM model response is a sequence of API calls, with each call consisting of an API name and its argument-value pairs. We use the Partial Sequence Matching metric to determine how many predicted APIs (with their argument-value pairs) in a sequence match with the gold API sequence. In contrast, the Full Sequence Matching metric evaluates whether the model predicts the exact full sequence of APIs, including both the API names and their argument-value pairs, when compared to the gold API sequence. In both cases, we calculate the scores for each sample and then compute the statistical mean across the entire dataset as the final score. We also use is the Win Rate, which measures if all the predicted APIs by the model are valid and when executed lead to the gold answer. Unlike the F1 and Accuracy measures, which assess the alignment of predictions with the gold API sequence, the Win Rate focuses on the final gold answer. A win is recorded if the predicted answer matches the gold answer.
@@ 4.4 Results
Table 1 presents a comparison of different baselines on the NESTful dataset with one-shot and three-shot ICL example settings. The low numbers of the best function calling models depict the complexity and toughness of the nested sequencing problem. GPT-4o and DeepSeek-V3 achieve the highest win-rate of 60%, which is significantly below the acceptable numbers for real-world applications in general. This clearly depicts the significant scope for improvements for the models in various aspects of function calling, including nested sequencing. We inspected the models’ outputs and identified several common issues across them. These models struggle with tasks such as assigning variables, utilizing output parameter details from the API specifications, and correctly passing variable names and corresponding output parameters to subsequent APIs, even with ICL examples [^5] Note: While we acknowledge that these models were not trained using the robust data schema outlined in Section 3.1, the challenges associated with nested sequencing persist regardless of the schema used and remain an area where LLMs need improvement..
As anticipated, in most of our experiments in Table 1, the models are doing better across all the metrics when they are provided with three ICL examples in the prompt instead of one example. Across all models, Partial Sequence Match scores are consistently higher than Full Sequence Match scores, which is expected, as the latter is a stricter metric than the former. In many cases, the Win Rate is higher than the Full Accuracy because models may follow alternative reasoning paths/trajectories to arrive at the correct final answer. While such deviations can penalize full or partial accuracy scores, they are still credited under the win rate for successfully reaching the gold answer. Hammer2.0-7b, despite being a smaller model, outperforms several larger tool-augmented LLMs. DeepSeek-V3 emerges as the strongest open-source model, closely matching GPT-4o’s performance in the three-shot ICL setting, although it trails slightly in the one-shot setup. These results highlight that model size or architectural complexity is not the primary determinant of performance; rather, the ability to effectively follow instructions and leverage in-context examples plays a more critical role. This is evident as some large models like xLAM-8x22b-fc-r and Llama-3-1-405B-Instruct-fp8 underperform, while smaller models like Hammer2.0-7b achieve exceptional results.
@@
Model Direct Prompting (One-shot ICL) ReAct Agent (Zero-shot)
Hammer2.0-7b 0.16 0.07
AgentLM-13B 0.00 0.00
Mixtral-8x22B-Instruct-v0.1 0.07 0.30
DeepSeek-V3 0.43 0.46
Table 2: Evaluation results (Win Rate) on NESTful comparing the performance of a ReAct Agent (zero-shot) to the Direct Prompting with a one-shot ICL example. For each model, the best performance is highlighted in bold for comparison.
Table 2 summarizes the results of the ReAct agent Yao et al. (2023) in compared to one-shot ICL direct prompting. Note that there was no ICL example provided in the ReAct case as the expected output does not need to follow the label assignment syntax. We only report Win Rate), which check if the trajectory of output tools lead to the gold answer, due to the single-step planning and execution approach of REACT as opposed to planning the entire API sequence at once in direct prompting.
For larger models like Mixtral-8x22B-Instruct-v0.1 and DeepSeek-V3, the ReAct approach outperforms direct prompting, though there is still room for improvement. Notably, Mixtral-8x22B-Instruct-v0.1 shows the highest win-rate gain of 30% with ReAct. Hammer2.0-7B performs better with direct prompting compared to the ReAct approach, Although AgentLM-13B is specifically trained for agentic tasks, it does not demonstrate strong performance on NESTful, indicating that agent-specific training or architectures do not always guarantee improved results in this setting. Output analysis reveals that the top-performing models exhibited better alignment with the ReAct format and occasionally relied on their parametric knowledge to replicate API functionalities. As a result, they achieved correct final outcomes, reflected in higher win-rate despite inconsistencies in intermediate steps.
@@ 4.5 Dataset Analysis
Figure 6: Top 4 models’ performances (on the three-shot setting) with a varying number of (a) longest data dependencies, (b) the total number of data dependencies, and (c) common data nesting patterns in NESTful. We observe that models perform well for the simple data samples, however, the performance drops sharply with complicated structural patterns in the data.
@@
In Figure 6(b), we present the win rate compared to the total number of data dependencies within a sample (a representation of the complexity of the sequence). The trends here are similar to the ones we observed in the Nesting Depth Analysis section. Models perform well for singular nested data dependency in the sample, achieving around 75% win-rate, but the performance drops sharply with two or more nested data dependencies.
@@ 5 Challenges
Data-type Adherence for the Input/Output Parameters   In the API specification, we define the data type for all parameters. The ‘type’ field specifies the data type, such as string, number, list, etc. Since APIs follow a strict structure for both input and output, it is crucial for the model to adhere to these specified formats. If the model fails to follow this, especially in cases involving nested functions where the output of one API is used as the input for another, the process will fail if the output type doesn’t match the expected input type.
@@
The example highlights the complexity of distinguishing repeated functions with different outputs, which models struggle with due to a lack of schema training—a challenge also supported by our qualitative analysis in Section 4.4.
Implicit API calling   Implicit function calling occurs when a model must identify and invoke the appropriate APIs to solve a user query, even though the query doesn’t explicitly mention them. This requires understanding the problem, selecting the correct functions, and filling in parameters using query details or previous outputs—adding significant complexity to the task. Figure 1 demonstrates an example of implicit function calling, where the user query presents an arithmetic problem without explicitly stating the APIs involved.
@@ 6 Conclusion
In this work, we introduced NESTful, a new benchmark for evaluating the LLMs on nested sequences of API function calling. Existing LLMs perform poorly on this dataset than on existing benchmarks. We also studied their performance and identified several modes of failure. In addition, we outlined many challenges this dataset poses to LLM function calling approaches. By releasing this dataset under a permissive open-source license, we aim to advance tool calling and enable solutions to more realistic, challenging tasks.